### v1

@@ Abstract
We provide a methodology for decoupling the bulk gravitational field equations of braneworld black holes to suppress the bulk singularities. To achieve this, we apply a linear Minimal Geometric Deformation (MGD) with respect to a coupling constant $\alpha$ to the $4D$ Minkowski spacetime embedded in an extra dimension. This results in a gravitational decoupling into a system $\mathcal{A}$ with equations of motion of order $\alpha^{0}$ and a system $\mathcal{B}$, related to the so-called Quasi-Einstein equations of order $\alpha$. This methodology allows for the construction of a regular geometry everywhere. We outline the necessary constraints for eliminating singularities and provide a recipe for solving the equations of motion. Both the warp factor, the scalar field, and the potential obtained are smooth and free from Dirac delta singularities. A control parameter is introduced such that, in the limit $b\to 0$, the Randall-Sundrum (RS) setup is recovered, resulting in a transition from a thick brane to a thin brane. The asymptotic behavior of the curvature invariant $\lim_{y\to\pm\infty}R_{5D}(r,y)$ is positive near the de Sitter core (for small $r$), asymptotically negative for finite $r>r_{*}$, and asymptotically flat at the $4D$ boundary as $r\to\infty$. Although this work aims to suppress bulk singularities, it is expected that our methodology may be useful for future investigations related to the embedding of gravitational objects within other braneworld contexts.
@@
The recent detection of gravitational waves by LIGO and VIRGO, coupled with the first image of a supermassive black hole captured by the Event Horizon Telescope (EHT) Abbott et al. (2016, 2017); Akiyama et al. (2022), has provided experimental data on gravity in the universe. These experiments have allowed us to test gravitational theories beyond General Relativity. In this context, braneworld theories have been tested using data from these experiments in recent years Banerjee et al. (2021, 2022); Visinelli et al. (2018); Mishra et al. (2022). The Randall-Sundrum (RS) model Randall and Sundrum (1999a, b) offers a geometric explanation for why gravity is weaker than the other fundamental forces of nature.
@@
On the other hand, in recent years, the Gravitational Decoupling (GD) algorithm Ovalle (2017) has garnered significant attention for yielding new solutions of physical interest. To name just a few cases, this method has been used to generate rotating black holes Contreras et al. (2021), hairy black holes Ovalle et al. (2021), spherically symmetric black holes Ovalle et al. (2018), regular black holes Ovalle et al. (2023), stellar distributions Estrada and Tello-Ortiz (2018); Estrada and Prado (2019), Charged Anisotropic Spherical Solutions Sharif and Sadiq (2018), solutions in Lovelock gravity Estrada (2019), among others. This method was also studied for braneworld models in reference León and Sotomayor (2019), but using the aforementioned methodology, based on the induced equations of motion on the brane. Broadly speaking, this method can be explained as follows: Initially, we have a gravitational system $\mathcal{A}$ whose geometry and matter source are of order $\alpha^{0}$. This geometry is linearly deformed with respect to a dimensionless coupling constant $\alpha$. This geometric deformation is coupled to a matter source, also of order $\alpha$, giving rise to a gravitational system $\mathcal{B}$, whose equations of motion are of order $\alpha$ and are so-called as Quasi-Einstein equations. Consequently, the equations of motion are decoupled into equations of order $\alpha^{0}$ for $\mathcal{A}$ and equations of order $\alpha$ for $\mathcal{B}$. First, the Einstein equations are solved for $\mathcal{A}$, followed by a simpler set of motion equations for $\mathcal{B}$. Finally, the two solutions are combined to derive the complete solution for the entire system.
@@ II.6.3 Ricci Scalar and regularity of the geometry :
Although the issue of stability under perturbations is beyond the scope of this work and could be explored in the future, it is worth mentioning that the behavior where $R_{5D}$ asymptotically approaches zero in the extra coordinate has been studied in Reference Germán et al. (2016) to address the stability problem associated with scalar perturbations of tachyonic thick branes. See other scenarios where $\lim_{y\to\pm\infty}R_{5D}\to 0$ in Koley and Kar (2005).
@@ III Discussion and Summary
To establish our methodology, we use the Minkowski spacetime embedded in an extra dimension as the $4D$ seed geometry, which is linearly deformed with respect to a coupling constant $\alpha$, referred to as Minimal Geometric Deformation (MGD). Thus, the gravitational system decouples in the following manner: a system $\mathcal{A}$ of order $\alpha^{0}$, which is coupled to the scalar field and potential matter sources. Conversely, the MGD of Minkowski space leads to the system $\mathcal{B}$, related with the so-called Quasi-Einstein equations, which are of order $\alpha$ and are coupled to the bulk energy-momentum tensor fields. Additionally, we have determined the properties that the bulk energy-momentum tensor must satisfy. It is important to note that both gravitational sectors $\mathcal{A}$ and $\mathcal{B}$ do not exchange energy and momentum; rather, there is only gravitational interaction between them. We will outline the constraints that must be satisfied to eliminate the singularity present both at the origin of the radial coordinate and at infinity in the extra coordinate. Additionally, we will provide a recipe for solving our equations of motion. Consequently, the equations of order $\alpha$ eliminate the singularity at the radial origin, while the equations of order $\alpha^{0}$ eliminate the singularity at infinity in the extra dimension.

### v2

@@ Abstract
We provide a methodology for decoupling the bulk gravitational field equations of braneworld black holes to suppress the bulk singularities. Thus, we provide a regular braneworld black hole setup. To achieve this, we apply a Minimal Geometric Deformation (MGD) with respect to a coupling constant $\alpha$ to the $4D$ Minkowski spacetime embedded in an extra dimension. This results in a gravitational decoupling into a system $\mathcal{A}$ with equations of motion of order $\alpha^{0}$ and a system $\mathcal{B}$, related to the so-called Quasi-Einstein equations of order $\alpha$. This methodology allows for the construction of a regular geometry everywhere. We outline the necessary constraints for eliminating singularities and provide a recipe for solving the equations of motion. Both the warp factor, the scalar field, and the potential obtained are smooth and free from Dirac delta singularities. A control parameter is introduced such that, in the limit $b\to 0$, the Randall-Sundrum (RS) setup is recovered, resulting in a transition from a thick brane to a thin brane. The asymptotic behavior of the curvature invariant $\lim_{y\to\pm\infty}R_{5D}(r,y)$ is positive near the de Sitter core (for small $r$), asymptotically negative for finite $r>r_{*}$, and asymptotically flat at the $4D$ boundary as $r\to\infty$. Although this work aims to suppress bulk singularities, it is expected that our methodology may be useful for future investigations related to the embedding of gravitational objects within other braneworld contexts.
@@
The recent detection of gravitational waves by LIGO and VIRGO, coupled with the first image of a supermassive black hole captured by the Event Horizon Telescope (EHT) Abbott et al. (2016, 2017); Akiyama et al. (2022), has provided experimental data on gravity in the universe. These experiments have allowed us to test gravitational theories beyond General Relativity. In this context, braneworld theories have been tested using data from these experiments in recent years Banerjee et al. (2021, 2022); Visinelli et al. (2018); Mishra et al. (2022). However, it is worth mentioning that the analyses conducted in these recent references seem not to consider the LHC results. Regarding this, some efforts to compare brane-world models with LHC results have been made in the literature. In reference Casadio and Ovalle (2012), the minimal geometric deformation of stars in the brane-world is studied, and brane-world black hole metrics are found. It is stated that a minimum mass for semiclassical microscopic black holes can be derived, which has a relevant impact on the description of black hole events at the LHC. Also, see reference da Rocha and Coimbra-Araujo (2006), where the possible effects of the brane world are investigated for mini-black holes and their measurable consequences in Large Hadron Collider observations. It is stated that if the fundamental Planck scale is at the TeV scale, the LHC will produce more than $10^{7}$ black holes per year. Therefore, contrasting brane-world models with experimental results seems to be an open issue. See more recent studies for example in references Das et al. (2020); Belis et al. (2024) The Randall-Sundrum (RS) model Randall and Sundrum (1999a, b) offers a geometric explanation for why gravity is weaker than the other fundamental forces of nature.
@@
On the other hand, in recent years, the Gravitational Decoupling (GD) algorithm Ovalle (2017) has garnered significant attention for yielding new solutions of physical interest. To name just a few cases, this method has been used to generate rotating black holes Contreras et al. (2021), hairy black holes Ovalle et al. (2021), spherically symmetric black holes Ovalle et al. (2018a), regular black holes Ovalle et al. (2023), stellar distributions Estrada and Tello-Ortiz (2018); Estrada and Prado (2019), Charged Anisotropic Spherical Solutions Sharif and Sadiq (2018), solutions in Lovelock gravity Estrada (2019), among others. This method was also studied for braneworld models in reference León and Sotomayor (2019), but using the aforementioned methodology, based on the induced equations of motion on the brane. Broadly speaking, this method is explained in reference Ovalle et al. (2018b) as follows: “given two gravitational sources $\mathcal{A}$ and $\mathcal{B}$, standard Einstein’s equations are first solved for $\mathcal{A}$, and then a simpler set of quasi-Einstein equations are solved for $\mathcal{B}$. Finally, the two solutions can be combined in order to derive the complete solution for the total system $\mathcal{A}\cup\mathcal{B}$.” It is worth mentioning that this method is not perturbative Sharif and Saba (2020); da Rocha (2022), but rather based on an exact gravitational decoupling.
@@ II.6.3 Ricci Scalar and regularity of the geometry :
On the other hand, although the stability under wavelength perturbations is beyond the scope of this work, the particular behavior of the Ricci curvature invariant, which in our case depends on both the radial coordinate and the extra dimension, raises the question of how our geometry would behave under such perturbations. In this context, regarding the specific case where $\lim_{y\to\pm\infty}R_{5D}\to 0$, references Gregory and Laflamme (1993); Gregory (2000) assert that black hole solutions embedded in an extra dimension within a braneworld framework (the so-called black string) can be unstable to wavelength perturbations in asymptotically flat space. This type of asymptote is also studied in Germán et al. (2016) to address the stability problem associated with scalar perturbations of tachyonic thick branes. Other scenarios where $\lim_{y\to\pm\infty}R_{5D}\to 0$ are discussed in Koley and Kar (2005). Thus, the stability of this type of asymptote, as well as the others previously described, could be investigated in future work.
II.7 A Brief Discussion on the Extension of the Horizon in the Bulk
We can easily detect the existence of an event horizon since the induced metric for a fixed value of $y$ in Equation (2) precisely represents a four-dimensional regular black hole. However, a natural question that arises when studying braneworld black holes is whether the event horizon of the black hole geometry extends into the bulk. This problem has been addressed in different contexts for the case of thin branes confined by a Dirac delta. The presence of the latter leads to a specific structure of the induced Einstein equations on the brane Shiromizu et al. (2000). Using this fact, in reference Casadio and Mazzacurati (2003), a method is proposed to determine the shape of the event horizon extension in the bulk. This method is applied for the case where, using the mentioned Dirac delta in the induced equations, the projected Ricci scalar on the brane vanishes, $R^{4D}=0$, and where the bulk equations in the directions of the brane are such that $R^{5D}_{ij}=\Lambda g_{ij}$ with $i,j=0,1,2,3$. Astrophysical data of the order of the solar mass were used. The analysis suggests that the horizon does close in the bulk, and a value for the extra coordinate $y=y_{0}$ is obtained, which marks the extent to which the horizon would propagate in the form of a pancake. In this same framework, reference Chamblin et al. (2001) argues that, in a braneworld black hole where the induced metric corresponds to a charged black hole, the event horizon extends into the dimension transversal to the brane depending on the value of the electric charge parameter.
However, it seems that the methodologies described in the references from the previous paragraph should not be applicable to our framework, due to the presence of a scalar field instead of a Dirac delta and the presence of matter fields in the bulk. Thus, a more in-depth study is needed in future work to explore the possible extension of the event horizon for thick branes in the presence of a scalar field.
II.7.1 Event horizon in the representation of an exponentially localized five-dimensional brane-world black hole.
Recent work in references Nakas and Kanti (2021a, b); Nakas et al. (2024) developed an algorithm that allows for the representation of the geometry of an analytic, exponentially localized five-dimensional brane-world black hole. This approach contrasts with the usual $4D$ black hole embedded in an extra dimension. Although the two representations differ, remarkably, this algorithm enables testing of how the event horizon extends in the bulk under this different representation. Therefore, although our previously described background is devoted to the usual thick brane embedded in an extra dimension, the use of this algorithm, based on the representation of an exponentially localized five-dimensional brane-world black hole, will allow us to test whether it is possible to investigate if the event horizon of out thick brane extends in the bulk under this representation. Thus, although the extension of the event horizon in this representation does not suggest that the event horizon would also extend in the usual background previously described, in this section we present an example of a thick brane-world black hole configuration where the extension of the horizon in the bulk is tested.
  1. 1.
    Consider the typical line element of RS:
    $ds^{2}=e^{2A(y)}\left(-dt^{2}+dr^{2}+r^{2}\,d\Omega_{2}^{2}\right)+dy^{2}$ (40)
  2. 2.
    Apply the invertible coordinate transformation $dy^{2}=e^{2A(y)}\,dz^{2}$ and we are led to
[…]
    In our case the function $A(y)$ is given by the equation (26). Thus, we obtain:
    $ds^{2}=e^{2A(y(z))}\left(-dt^{2}+dr^{2}+r^{2}\,d\Omega_{2}^{2}\right)+dz^{2}$ (42)
    where $d\Omega_{2}^{2}=d\theta^{2}+\sin^{2}\theta\,d\phi^{2}$.
  3. 3.
    Next, we rewrite the flat component of the line element (40) in five-dimensional spherical coordinates by using the following transformation:
    $\{r,z\}\to\{\rho\sin\chi,\,\rho\cos\chi\},\quad\chi\in[0,\pi]$ (43)
    where, the inverse transformation is given by:
    $\{\rho,\chi\}\to\left\{\sqrt{r^{2}+z^{2}},\,\tan^{-1}\left(\frac{r}{z}\right)\right\}$ (44)
    Thus, the equation (42) is:
    $ds^{2}=e^{2A(y(\rho\cos\chi))}\left(-\,dt^{2}+d\rho^{2}+\rho^{2}\,d\Omega_{3}^{2}\right)$ (45)
    where $d\Omega_{3}^{2}=d\chi^{2}+\sin^{2}\chi\,d\theta^{2}+\sin^{2}\chi\sin^{2}\theta\,d\phi^{2}$.
  4. 4.
    It is proposed to rewrite the last equation as follows:
    $ds^{2}=e^{2A(y(\rho\cos\chi))}\left(-U(\rho)dt^{2}+\frac{d\rho^{2}}{U(\rho)}+\rho^{2}\,d\Omega_{3}^{2}\right)$ (46)
    It is worth mentioning that the above line element represents an exponentially localized five-dimensional brane-world black hole. See figure 6.
  5. 5.
    It is possible to express the line element (46) in the original coordinate system $\{t,r,\theta,\phi,y\}$ by following the inverse transformations (44):
    $ds^{2}=e^{2A(y)}\Big{[}-U(r,y)\,dt^{2}+\frac{dr^{2}}{r^{2}+z^{2}(y)}\left(r^{2}/U(r,y)+z^{2}(y)\right)+r^{2}d\Omega^{2}_{2}$
    $+\frac{2rz(y)e^{A(y)}}{r^{2}+z^{2}(y)}\left(1/U(r,y)-1\right)\,dr\,dy\Big{]}$
    $+\frac{dy^{2}}{r^{2}+z^{2}(y)}\left(r^{2}+z^{2}(y)/U(r,y)\right)$ (47)
    We can see that the induced metric at $y,z\to 0$ is
    $ds^{2}=e^{2A(0)}\left(-U(r)dt^{2}+U(r)^{-1}dr^{2}+r^{2}d\Omega_{2}^{2}\right)$ (48)
    where $e^{2A(0)}$ is given by the equation (32).
  6. 6.
    In reference Nakas and Kanti (2021a) is defined $U(r,y_{0})=1-\bar{m}/\sqrt{r^{2}+z(y_{0})^{2}}$. As a test for this subsection, we will use a generalization of Hayward’s mass function (since Dymnikova’s mass function grows exponentially and very rapidly). In order to be consistent with the algorithm, we will define the $U$ function as:
    $U(r,y_{0})=1-\alpha\cdot\frac{m(r,y_{0})}{\sqrt{r^{2}+z(y_{0})^{2}}}=1-\alpha\cdot\bar{m}\frac{(\sqrt{r^{2}+z(y_{0})^{2}}))^{3}}{\left((\sqrt{r^{2}+z(y_{0})^{2}})^{3}+\bar{m}\right)\sqrt{r^{2}+z(y_{0})^{2}}}$ (49)
    where $\bar{m}$ is a constant and $z(y_{0})$ is given by equation (41) for $y=y_{0}$. We define the mass parameter to the value of $\bar{m}$ in the parameter space such that $U(r_{0},y_{0})=0$.
    $\bar{m}=\frac{(r_{0}^{2}+z(y_{0})^{2})\sqrt{r_{0}^{2}+z(y_{0})^{2}}}{-1+2r_{0}^{2}+2z(y_{0})^{2}}$ (50)
    Thus, for a value of $\bar{m}$, there exist one or more ordered pairs $\left(\bar{m},(r_{0},y_{0})\right)$ such that $U(r_{0},y_{0})=0$. The smallest value of $r_{0}$ represents the inner horizon, and the largest value represents the event horizon. Since $z(y)$ cannot be computed analytically, in Figures 7, we compute the value of $\bar{m}$ numerically We can intuitively notice in the figure above that there are two ordered pairs $(r_{0}=r_{i},y_{0}=y_{i})$ and $(r_{0}=r_{h},y_{0}=y_{h})$ such that $r_{i}<r_{h}$ and $y_{i}<y_{h}$, and both ordered pairs correspond to the same value of the mass parameter $\bar{m}$. In this way, we can intuit that the inner horizon would be located at $(r_{0}=r_{i},y_{0}=y_{i})$ and the event horizon would be located at $(r_{0}=r_{h},y_{0}=y_{h})$. Furthermore, in the graph below, $\sqrt{r_{0}^{2}+z(y_{0})^{2}}$ is plotted on the horizontal axis and $\bar{m}$ on the vertical axis. From this graph, we can infer that there is a critical value $\bar{m}=\bar{m}_{c}$ such that for $\bar{m}<\bar{m}_{c}$ there is no black hole, for $\bar{m}=\bar{m}_{c}$ we have an extremal black hole, and for $\bar{m}>\bar{m}_{c}$, we have a black hole with an inner horizon (located on the curve with negative slope) and an event horizon (located on the curve with positive slope). Thus, we can intuit that both the inner horizon $r_{i}$ and the event horizon $r_{h}$ may extend into the bulk for values $y_{i}$ and $y_{h}$, respectively, in the parameter space. Unfortunately, since our analysis is numerical, it is difficult to estimate whether the horizon extends in the bulk in the form of a pancake as in Nakas and Kanti (2021a) or another shape, and whether there is a value in the bulk at which the event horizon ceases to exist It is worth mentioning that the representation of an exponentially localized five-dimensional brane-world black hole also implies the variation of other physical features, such as energy conditions, thermodynamics, etc. These aspects go beyond the scope of this work, as in this paper, we have only provided an overview of the potential extension of the event horizon in the bulk.
Figure 6: brane world set up
Figure 7: $\bar{m}$ parameter for $k=1,b=0.4,c=100$
@@ III Discussion and Summary
To establish our methodology, we use the Minkowski spacetime embedded in an extra dimension as the $4D$ seed geometry, to which a Minimal Geometric Deformation (MGD) is applied. Thus, the gravitational system decouples in the following manner: a system $\mathcal{A}$ of order $\alpha^{0}$, which is coupled to the scalar field and potential matter sources. Conversely, the MGD of Minkowski space leads to the system $\mathcal{B}$, related with the so-called Quasi-Einstein equations, which are of order $\alpha$ and are coupled to the bulk energy-momentum tensor fields. Additionally, we have determined the properties that the bulk energy-momentum tensor must satisfy. It is important to note that both gravitational sectors $\mathcal{A}$ and $\mathcal{B}$ do not exchange energy and momentum; rather, there is only gravitational interaction between them. We will outline the constraints that must be satisfied to eliminate the singularity present both at the origin of the radial coordinate and at infinity in the extra coordinate. Additionally, we will provide a recipe for solving our equations of motion. Consequently, the equations of order $\alpha$ eliminate the singularity at the radial origin, while the equations of order $\alpha^{0}$ eliminate the singularity at infinity in the extra dimension.